Determine which of the following is :
a) the most concentrated, and
b) the most dilute.
(i) 0.45 mol L-1 NaCI (ii) 4.2 % m/v NaCI
(iii) 41 .6 g NaCI L-1 (iv) 39950 ppm NaCI
1
Expert's answer
2020-07-30T07:10:48-0400

Let convert all to mol L-1.

(ii): n(NaCl) = 4.2g/58.5=0.0718 mol. c(NaCl)= 0.0718/0.1=0.718 mol L-1

(iii): n(NaCl) = 41.6g/58.5=0.711 mol.

c(NaCl)= 0.711/1=0.711 mol L-1

(iv): n(NaCl) = 39.95g/58.5=0.683 mol.

c(NaCl)= 0.683/1=0.683 mol L-1

Thus the most concentrated - (ii)

the most dilute - (i)